# Possible Existence of $^{3}_{\phi}$H, $^{4}_{\phi}$H, $^{4}_{\phi}$He, and $^{5}_{\phi}$He Nuclei

### v1

@@ Abstract
In this Letter, we report the first systematic few-body study of $\phi$-mesic systems up to five particles. Our approach integrates HAL QCD–derived $\phi N$ interactions [23, 24] combined with the MT I–III $NN$ potential [31, 32]. Faddeev–Yakubovsky equations [33] for two distinct particle species were formulated in configuration-space and solved numerically, extending earlier four- and five-body frameworks [34, 35, 36, 37]. This approach enables the first systematic and unified analysis of $\phi NNN$ and $\phi NNNN$ bound states, constrained by lattice QCD dynamics.
@@
For the nucleon–nucleon interaction, we employ the MT I–III potential [31, 32], $V_{NN}^{S}(r)=v_{r}\frac{e^{-\mu_{r}r}}{r}+v_{a}^{s}\frac{e^{-\mu_{a}r}}{r},$ with $\mu_{r}=3.11$ fm-1, $\mu_{a}=1.55$ fm-1, and $v_{r}=1438.72$ MeV·fm. Standard strengths are $v_{a}^{s=0}=-513.968$ and $v_{a}^{s=1}=-626.885$ MeV·fm, which allow for reproducing fairly well the deuteron binding energy and $np$ scattering lengths using physical masses. For calculations employing QCD-motivated nucleon masses we readjust to $v_{a}^{s=0}=-510.493$ MeV·fm and $v_{a}^{s=1}=-623.269$ MeV·fm, yielding $B_{d}=2.22457$ MeV and $a_{np}=23.737$ fm.
@@
Using the isospin formalism, neutrons and protons can be treated as identical nucleons that differ only in the projection of their isospin quantum number. This particle-permutation symmetry reduces the number of independent FYE components and, consequently, the number of coupled differential equations to solve. In this work, we consider $\phi$-mesic systems: systems of $A$-particles containing one particle that differs from the remaining A1A-1 identical nucleons.
@@
Thus, one can reduce the number of independent FYE components to 2, 5, and 16, when solving $A=3$, 4 and 5 particle problems, respectively. The corresponding numbers of independent equations are summarized in Table 1. We emphasize that the developed theoretical framework enables rigorous few-body calculations of light hypernuclei containing three or four nucleons bound to baryons such as $\Lambda$, $\Omega$, and $\Xi$. In particular, it allows for a systematic investigation of the $YNNN$ and YNNNNYNNNN ($Y=\Lambda$, $\Omega$, $\Omega_{ccc}$, and $\Xi$) systems within a fully microscopic few-body approach. In the case of $YNNNN$ systems, the applicability of the $Y\alpha$ cluster models can be assessed through direct five-particle calculations.
@@ Acknowledgments
This work is supported by the City University of New York PSC CUNY Research Award No. 68541-00 915 56, and the Department of Energy/National Nuclear Security Administration under Award No. NA0003979 and by French IN2P3 for a theory project “PUMA”. We were granted access to the HPC resources of TGCC/IDRIS under the allocation 2025-AD010506006R3 made by GENCI (Grand Equipement National de Calcul Intensif).

### v2

@@ Abstract
In this Letter, we report the first systematic few-body study of $\phi$-mesic systems up to five particles. Our approach integrates HAL QCD–derived $\phi N$ interactions [23, 24] combined with the MT I–III NNNN potential [31, 32]. Faddeev–Yakubovsky equations [33] for two distinct particle species were formulated in configuration-space and solved numerically, extending earlier four- and five-body frameworks [34, 35, 36, 37]. This approach enables the first systematic and unified analysis of $\phi NNN$ and $\phi NNNN$ bound states, constrained by lattice QCD dynamics.
@@
For the nucleon–nucleon interaction, we employ the MT I–III potential [31, 32], $V_{NN}^{s}(r)=v_{r}\frac{e^{-\mu_{r}r}}{r}+v_{a}^{s}\frac{e^{-\mu_{a}r}}{r},$ with $\mu_{r}=3.11$ fm-1, $\mu_{a}=1.55$ fm-1, and $v_{r}=1438.72$ MeV·fm. Standard strengths are $v_{a}^{s=0}=-513.968$ and $v_{a}^{s=1}=-626.885$ MeV·fm, which allow for reproducing fairly well the deuteron binding energy and npnp scattering lengths using physical masses. For calculations employing QCD-motivated nucleon masses we readjust to $v_{a}^{s=0}=-510.493$ MeV·fm and $v_{a}^{s=1}=-623.269$ MeV·fm, yielding $B_{d}=2.2246$ MeV and $a_{np}=23.737$ fm.
@@
Using the isospin formalism, neutrons and protons can be treated as identical nucleons that differ only in the projection of their isospin quantum number. This particle-permutation symmetry reduces the number of independent FYE components and, consequently, the number of coupled differential equations to solve. In this work, we consider $\phi$-mesic systems: systems of AA-particles containing one particle that differs from the remaining A1A-1 identical nucleons.
@@
Thus, one can reduce the number of independent FYE components to 2, 5, and 16, when solving $A=3$, 4 and 5 particle problems, respectively. The corresponding numbers of independent equations are summarized in Table 1. We emphasize that the developed theoretical framework enables rigorous few-body calculations of light hypernuclei containing three or four nucleons bound to baryons such as $\Lambda$, $\Omega$, and $\Xi$. In particular, it allows for a systematic investigation of the YNNNYNNN and YNNNNYNNNN ($Y=\Lambda$, $\Omega$, $\Omega_{ccc}$, and $\Xi$) systems within a fully microscopic few-body approach. In the case of YNNNNYNNNN systems, the applicability of the $Y\alpha$ cluster models can be assessed through direct five-particle calculations.
@@ Acknowledgments
This work is supported by the City University of New York PSC CUNY Research Award No. 68541-$0$ 915 56, and the Department of Energy/National Nuclear Security Administration under Award No. NA0003979 and by French IN2P3 for a theory project “PUMA”. We were granted access to the HPC resources of TGCC/IDRIS under the allocation 2025-AD010506006R3 made by GENCI (Grand Equipement National de Calcul Intensif).